OpenStudy (anonymous):

Can someone help me step by step? derivative y=ln(sqrt(x))+sec^-1(x) I know arcsec is 1/xsqrt(x^2-1)

OpenStudy (anonymous):

ln(x^1/2)= 1/2*ln(x) does that help

OpenStudy (anonymous):

how does 1/2ln(x) turn into 1/2x?

OpenStudy (anonymous):

ln(x) = 1/x from the table

OpenStudy (anonymous):

1/2*1/x=1/2x

OpenStudy (anonymous):

Oh I didn't notice that. Thank you!

OpenStudy (anonymous):

no problem